OpenStudy (anonymous):

Let g(x) be the reflection of f(x) = 3x + 2 in the x-axis. What is a function rule for g(x)?

OpenStudy (anonymous):

g(x) = 3x – 2 g(x) = –3x – 2 g(x) = 3x + 2 g(x) = –3x + 2

OpenStudy (anonymous):

When you reflect a function in the x-axis, the x coordinates stay the same and the y coordinates are replaced by -y. i.e., f(x) becomes g(x)=-f(x) after reflecting in the x axis.